# Soft Theorems in Chern-Simons Matter Theories

### v3

@@ 2 Subleading Soft Theorem from Sen’s Covariantisation Procedure
$\,\,=\frac{i}{2}\left[\frac{\partial K^{\alpha\gamma}}{\partial p_{\mu}}\left(-p-k\right)e_{\mu}{}_{\gamma}^{\,\,\,\,\beta}-\frac{1}{2}\frac{\partial^{2}K^{\alpha\gamma}}{\partial p_{\mu}\partial p_{\nu}}\left(-p-k\right)k_{\mu}e_{\nu}{}_{\gamma}^{\,\,\,\,\beta}-\frac{\partial K^{\beta\gamma}}{\partial p_{\mu}}\left(p\right)e_{\mu}{}_{\gamma}^{\,\,\,\,\alpha}\right$.
@@
$\,\,=\frac{i}{2}\left[\frac{\partial K\left(-p-k\right)}{\partial p_{\mu}}e_{\mu}-\frac{1}{2}\frac{\partial^{2}K\left(-p-k\right)}{\partial p_{\mu}\partial p_{\nu}}k_{\mu}e_{\nu}-\left(\frac{\partial K\left(p\right)}{\partial p_{\mu}}e_{\mu}\right)^{T}-\frac{1}{2}\left(\frac{\partial^{2}K\left(p\right)}{\partial p_{\mu}\partial p_{\nu}}k_{\mu}e_{\nu}\right)^{T}\right]$
$\,\,=\frac{i}{2}\left[\frac{\partial K\left(-p\right)}{\partial p_{\mu}}e_{\mu}+\frac{1}{2}\frac{\partial^{2}K\left(-p\right)}{\partial p_{\mu}\partial p_{\nu}}k_{\mu}e_{\nu}-\left(\frac{\partial K\left(p\right)}{\partial p_{\mu}}e_{\mu}\right)^{T}-\frac{1}{2}\left(\frac{\partial^{2}K\left(p\right)}{\partial p_{\mu}\partial p_{\nu}}k_{\mu}e_{\nu}\right)^{T}\right]$
$\,\,=\frac{i}{2}\left[\frac{\partial K\left(-p\right)}{\partial p_{\mu}}e_{\mu}+\frac{1}{2}\frac{\partial^{2}K\left(-p\right)}{\partial p_{\mu}\partial p_{\nu}}k_{\mu}e_{\nu}-e_{\mu}^{T}\frac{\partial K^{T}\left(-p\right)}{\partial p_{\mu}}-\frac{1}{2}k_{\mu}e_{\nu}^{T}\frac{\partial^{2}K^{T}\left(-p\right)}{\partial p_{\mu}\partial p_{\nu}}\right]$
$\,\,=i\left[\frac{\partial K\left(-p\right)}{\partial p_{\mu}}e_{\mu}+\frac{1}{2}\frac{\partial^{2}K\left(-p\right)}{\partial p_{\mu}\partial p_{\nu}}k_{\mu}e_{\nu}\right],$ (19)
where I’ve used (8) in the second last line of (19) and (11) in the last line of (19). If the soft boson line in $\Gamma^{\left(3\right)}$ is an internal line, then (19) needs to be modified by stripping of the polarization vector and contracting the index with the appropriate soft propagator. Using (19) and the relations (11), (13), (15) and (16), we can write-
@@
$\,\,=i\left[\frac{\partial K\left(-p\right)}{\partial p_{\mu}}e_{\mu}+\frac{1}{2}\frac{\partial^{2}K\left(-p\right)}{\partial p_{\mu}\partial p_{\nu}}k_{\mu}e_{\nu}\right]\Xi\left(-p-k\right)$
$\,\,=-i\left[\frac{\partial K\left(-p\right)}{\partial p_{\mu}}\Xi\left(-p-k\right)+\frac{1}{2}\frac{\partial^{2}K\left(-p\right)}{\partial p_{\mu}\partial p_{\nu}}\Xi\left(-p-k\right)k_{\nu}\right]e_{\mu}^{T}$
$\,\,=-i\left[2ip^{\mu}-K\left(-p\right)\frac{\partial\Xi\left(-p\right)}{\partial p_{\mu}}+\frac{\partial K\left(-p\right)}{\partial p_{\mu}}\frac{\partial\Xi\left(-p\right)}{\partial p_{\nu}}k_{\nu}+\frac{1}{2}\frac{\partial^{2}K\left(-p\right)}{\partial p_{\mu}\partial p_{\nu}}\Xi\left(-p\right)k_{\nu}\right]e_{\mu}^{T}$
$\,\,=2p^{\mu}e_{\mu}^{T}+iK\left(-p\right)\frac{\partial\Xi\left(-p\right)}{\partial p_{\mu}}e_{\mu}^{T}+\frac{i}{2}\frac{\partial K\left(-p\right)}{\partial p_{\mu}}\frac{\partial\Xi\left(-p\right)}{\partial p_{\nu}}\left(k_{\mu}e_{\nu}-k_{\nu}e_{\mu}\right)^{T}$
$\,\,\,\,\,\,\,+\frac{i}{2}K\left(-p\right)\frac{\partial^{2}\Xi\left(-p\right)}{\partial p_{\mu}\partial p_{\nu}}k_{\nu}e_{\mu}^{T}.$ (20)
@@
$\left.+\frac{i}{4}\varepsilon_{i;\alpha}\frac{\partial K^{\alpha\gamma}\left(-p_{i}\right)}{\partial p_{i\nu}}\frac{\partial\Xi_{\gamma\delta}\left(-p_{i}\right)}{\partial p_{i\mu}}\left(e_{\mu}^{T}k_{\nu}-e_{\nu}^{T}k_{\mu}\right)_{\,\,\,\,\beta_{i}}^{\delta}\right]G_{n}^{\left(i\right);\beta_{i}}\left(p_{i}\right)$
$=$ $\frac{1}{p_{i}.k}\left[\varepsilon_{i}^{T}p_{i}^{\mu}e_{\mu}^{T}+k_{\rho}\varepsilon_{i}^{T}p_{i}^{\mu}e_{\mu}^{T}\frac{\partial}{\partial p_{i\rho}}+\frac{i}{4}\varepsilon_{i}^{T}\frac{\partial K_{i}\left(-p_{i}\right)}{\partial p_{i\nu}}\frac{\partial\Xi_{i}\left(-p_{i}\right)}{\partial p_{i\mu}}\left(k_{\nu}e_{\mu}-k_{\mu}e_{\nu}\right)^{T}\right]G_{n}^{\left(i\right)}\left(p_{i}\right),$ (21)
@@
Summing over soft boson attaching to all distinct external legs for $A1$ and adding to $A2$ we get,
@@
$+\sum_{i}\frac{1}{p_{i}.k}\frac{i}{4}\varepsilon_{i}^{T}\frac{\partial K_{i}\left(-p_{i}\right)}{\partial p_{i\nu}}\frac{\partial\Xi_{i}\left(-p_{i}\right)}{\partial p_{i\mu}}\left(k_{\nu}e_{\mu}-k_{\mu}e_{\nu}\right)^{T}G_{n}^{\left(i\right)}\left(p_{i}\right).$ (22)
Notice that the the first two terms in (22) are independent of $K$ and $\Xi$, whereas the last term depends on their derivatives and hence depends explicitly on the action.
@@
In the case of soft photons, the gauge group is simply $U\left(1\right)$ and the gauge field is abelian. There is only one generator for the lie algebra which is denoted by $Q_{\alpha}^{\,\,\,\,\beta}$ and is sometimes called the charge matrix since its matrix elements are charges of the various fields $\phi_{\alpha}$ in the external state. Therefore we have,
@@
$+\sum_{i}\frac{1}{p_{i}.k}\frac{i}{4}\varepsilon_{i}^{T}\frac{\partial K_{i}\left(-p_{i}\right)}{\partial p_{i\nu}}\frac{\partial\Xi_{i}\left(-p_{i}\right)}{\partial p_{i\mu}}Q^{T}\left(k_{\nu}e_{\mu}-k_{\mu}e_{\nu}\right)G_{n}^{\left(i\right)}\left(p_{i}\right).$ (23)
@@
$+\sum_{i}\frac{1}{p_{i}.k}\frac{i}{4}\varepsilon_{i}^{T}q_{i}\frac{\partial K_{i}\left(-p_{i}\right)}{\partial p_{i\nu}}\frac{\partial\Xi_{i}\left(-p_{i}\right)}{\partial p_{i\mu}}\left(k_{\nu}e_{\mu}-k_{\mu}e_{\nu}\right)G_{n}^{\left(i\right)}\left(p_{i}\right).$ (24)
@@
$\Xi_{i}\left(p_{i}\right)=i\left(\cancel{p_{i}}+m_{i}\right),$
@@
$\frac{i}{4}\left(e_{\mu}k_{\nu}-e_{\nu}k_{\mu}\right)q_{i}\varepsilon_{i}^{T}\frac{\partial K_{i}\left(-p_{i}\right)}{\partial p_{i\nu}}\frac{\partial\Xi_{i}\left(-p_{i}\right)}{\partial p_{i\mu}}=$ $-\frac{i}{4}q_{i}\varepsilon_{i}^{T}\left(e_{\mu}k_{\nu}-e_{\nu}k_{\mu}\right)i\gamma^{\nu}\gamma^{\mu}$
@@
$=$ $-iq_{i}\varepsilon_{i}^{T}\frac{k_{\mu}e_{\nu}}{p_{i}.k}S_{\left(i\right)}^{\mu\nu}.$
@@
$G_{n+1}\left(e,k;\varepsilon_{i},p_{i}\right)=\sum_{i}\varepsilon_{i}^{T}q_{i}\frac{e.p_{i}}{p_{i}.k}G_{n}^{\left(i\right)}\left(p_{i}\right)-i\sum_{i}\varepsilon_{i}^{T}q_{i}\frac{e_{\mu}k_{\nu}}{p_{i}.k}\left(\hat{J}_{\left(i\right)}^{\mu\nu}G_{n}^{\left(i\right)}\right)\left(p_{i}\right),$ (25)
@@
$\left(\hat{J}_{\left(i\right)}^{\mu\nu}G_{n}^{\left(i\right)}\right)\left(p_{i}\right)=i\left\{p_{i}^{\mu}\frac{\partial}{\partial p_{i\nu}}-p_{i}^{\nu}\frac{\partial}{\partial p_{i\mu}}\right\}G_{n}^{\left(i\right)}\left(p_{i}\right)+\frac{i}{4}\left[\gamma^{\mu},\gamma^{\nu}\right]G_{n}^{\left(i\right)}\left(p_{i}\right).$ (26)
@@
For gluons, the gauge group is $SU\left(3\right)$ and the Lie Algebra has 8 generators. If the external states are quarks, they will be in the fundamental representation of $SU\left(3\right)$ in which case the generators are given in terms of the 8 Gell-Mann matrices $\lambda_{a}$,
@@
where $\lambda_{i;a}$ are such that the $i$th block is the corresponding Gell-Mann matrix while the other blocks are identity. If the external states are gluons then the generators are in the adjoint representation of $SU\left(3\right)$,
@@
$G_{n+1}\left(e,k;\varepsilon_{i},p_{i}\right)=\sum_{i}\varepsilon_{i}^{T}g_{i}\frac{e_{a}.p_{i}}{p_{i}.k}T_{i;a}G_{n}^{\left(i\right)}\left(p_{i}\right)-i\sum_{i}\varepsilon_{i}^{T}g_{i}\frac{e_{a\mu}k_{\nu}}{p_{i}.k}T_{i;a}\left(\hat{J}_{\left(i\right)}^{\mu\nu}G_{n}^{\left(i\right)}\right)\left(p_{i}\right).$ (29)

### v4

@@ 2 Subleading Soft Theorem from Sen’s Covariantisation Procedure
$\,\,=\frac{\mathrm{i}}{2}\left[\frac{\partial K^{\alpha\gamma}}{\partial p_{\mu}}\left(-p-k\right)e_{\mu}{}_{\gamma}^{\,\,\,\,\beta}-\frac{1}{2}\frac{\partial^{2}K^{\alpha\gamma}}{\partial p_{\mu}\partial p_{\nu}}\left(-p-k\right)k_{\mu}e_{\nu}{}_{\gamma}^{\,\,\,\,\beta}-\frac{\partial K^{\beta\gamma}}{\partial p_{\mu}}\left(p\right)e_{\mu}{}_{\gamma}^{\,\,\,\,\alpha}\right$.
@@
$\,\,=\frac{\mathrm{i}}{2}\left[\frac{\partial K\left(-p-k\right)}{\partial p_{\mu}}e_{\mu}-\frac{1}{2}\frac{\partial^{2}K\left(-p-k\right)}{\partial p_{\mu}\partial p_{\nu}}k_{\mu}e_{\nu}-\left(\frac{\partial K\left(p\right)}{\partial p_{\mu}}e_{\mu}\right)^{T}-\frac{1}{2}\left(\frac{\partial^{2}K\left(p\right)}{\partial p_{\mu}\partial p_{\nu}}k_{\mu}e_{\nu}\right)^{T}\right]$
$\,\,=\frac{\mathrm{i}}{2}\left[\frac{\partial K\left(-p\right)}{\partial p_{\mu}}e_{\mu}+\frac{1}{2}\frac{\partial^{2}K\left(-p\right)}{\partial p_{\mu}\partial p_{\nu}}k_{\mu}e_{\nu}-\left(\frac{\partial K\left(p\right)}{\partial p_{\mu}}e_{\mu}\right)^{T}-\frac{1}{2}\left(\frac{\partial^{2}K\left(p\right)}{\partial p_{\mu}\partial p_{\nu}}k_{\mu}e_{\nu}\right)^{T}\right]$
$\,\,=\frac{\mathrm{i}}{2}\left[\frac{\partial K\left(-p\right)}{\partial p_{\mu}}e_{\mu}+\frac{1}{2}\frac{\partial^{2}K\left(-p\right)}{\partial p_{\mu}\partial p_{\nu}}k_{\mu}e_{\nu}-e_{\mu}^{T}\frac{\partial K^{T}\left(-p\right)}{\partial p_{\mu}}-\frac{1}{2}k_{\mu}e_{\nu}^{T}\frac{\partial^{2}K^{T}\left(-p\right)}{\partial p_{\mu}\partial p_{\nu}}\right]$
$\,\,=\mathrm{i}\left[\frac{\partial K\left(-p\right)}{\partial p_{\mu}}e_{\mu}+\frac{1}{2}\frac{\partial^{2}K\left(-p\right)}{\partial p_{\mu}\partial p_{\nu}}k_{\mu}e_{\nu}\right],$ (18)
where we’ve used (7) in the second last line of (18) and (10) in the last line of (18). If the soft boson line in $\Gamma^{\left(3\right)}$ is an internal line, then (18) needs to be modified by stripping of the polarization vector and contracting the index with the appropriate soft propagator. Using (18) and the relations (10), (12), (14) and (15), we can write-
@@
$\,\,=\mathrm{i}\left[\frac{\partial K\left(-p\right)}{\partial p_{\mu}}e_{\mu}+\frac{1}{2}\frac{\partial^{2}K\left(-p\right)}{\partial p_{\mu}\partial p_{\nu}}k_{\mu}e_{\nu}\right]\Xi\left(-p-k\right)$
$\,\,=-\mathrm{i}\left[\frac{\partial K\left(-p\right)}{\partial p_{\mu}}\Xi\left(-p-k\right)+\frac{1}{2}\frac{\partial^{2}K\left(-p\right)}{\partial p_{\mu}\partial p_{\nu}}\Xi\left(-p-k\right)k_{\nu}\right]e_{\mu}^{T}$
$\,\,=-\mathrm{i}\left[2\mathrm{i}p^{\mu}-K\left(-p\right)\frac{\partial\Xi\left(-p\right)}{\partial p_{\mu}}+\frac{\partial K\left(-p\right)}{\partial p_{\mu}}\frac{\partial\Xi\left(-p\right)}{\partial p_{\nu}}k_{\nu}+\frac{1}{2}\frac{\partial^{2}K\left(-p\right)}{\partial p_{\mu}\partial p_{\nu}}\Xi\left(-p\right)k_{\nu}\right]e_{\mu}^{T}$
$\,\,=2p^{\mu}e_{\mu}^{T}+\mathrm{i}K\left(-p\right)\frac{\partial\Xi\left(-p\right)}{\partial p_{\mu}}e_{\mu}^{T}+\frac{\mathrm{i}}{2}\frac{\partial K\left(-p\right)}{\partial p_{\mu}}\frac{\partial\Xi\left(-p\right)}{\partial p_{\nu}}\left(k_{\mu}e_{\nu}-k_{\nu}e_{\mu}\right)^{T}$
$\,\,\,\,\,\,\,+\frac{\mathrm{i}}{2}K\left(-p\right)\frac{\partial^{2}\Xi\left(-p\right)}{\partial p_{\mu}\partial p_{\nu}}k_{\nu}e_{\mu}^{T}.$ (19)
@@
$\left.+\frac{\mathrm{i}}{4}\varepsilon_{i;\alpha}\frac{\partial K^{\alpha\gamma}\left(-p_{i}\right)}{\partial p_{i\nu}}\frac{\partial\Xi_{\gamma\delta}\left(-p_{i}\right)}{\partial p_{i\mu}}\left(e_{\mu}^{T}k_{\nu}-e_{\nu}^{T}k_{\mu}\right)_{\,\,\,\,\beta_{i}}^{\delta}\right]G_{n}^{\left(i\right);\beta_{i}}\left(p_{i}\right)$
$=$ $\frac{1}{p_{i}.k}\left[\varepsilon_{i}^{T}p_{i}^{\mu}e_{\mu}^{T}+k_{\rho}\varepsilon_{i}^{T}p_{i}^{\mu}e_{\mu}^{T}\frac{\partial}{\partial p_{i\rho}}+\frac{\mathrm{i}}{4}\varepsilon_{i}^{T}\frac{\partial K_{i}\left(-p_{i}\right)}{\partial p_{i\nu}}\frac{\partial\Xi_{i}\left(-p_{i}\right)}{\partial p_{i\mu}}\left(k_{\nu}e_{\mu}-k_{\mu}e_{\nu}\right)^{T}\right]G_{n}^{\left(i\right)}\left(p_{i}\right),$ (20)
@@
Summing over the soft boson attaching to all distinct external legs for $A1$ and adding to $A2$ we get,
@@
$+\sum_{i}\frac{1}{p_{i}.k}\frac{\mathrm{i}}{4}\varepsilon_{i}^{T}\frac{\partial K_{i}\left(-p_{i}\right)}{\partial p_{i\nu}}\frac{\partial\Xi_{i}\left(-p_{i}\right)}{\partial p_{i\mu}}\left(k_{\nu}e_{\mu}-k_{\mu}e_{\nu}\right)^{T}G_{n}^{\left(i\right)}\left(p_{i}\right).$ (21)
Notice that the first two terms in (21) are independent of $K$ and $\Xi$, whereas the last term depends on their derivatives and hence depends explicitly on the action.
@@
In the case of soft photons, the gauge group is simply $\text{U}\left(1\right)$ and the gauge field is Abelian. There is only one generator for the lie algebra which is denoted by $Q_{\alpha}^{\,\,\,\,\beta}$ and is sometimes called the charge matrix since its matrix elements are the electric charges of the various fields $\phi_{\alpha}$ in the external state. Therefore we have,
@@
$+\sum_{i}\frac{1}{p_{i}.k}\frac{\mathrm{i}{}}{4}\varepsilon_{i}^{T}\frac{\partial K_{i}\left(-p_{i}\right)}{\partial p_{i\nu}}\frac{\partial\Xi_{i}\left(-p_{i}\right)}{\partial p_{i\mu}}Q^{T}\left(k_{\nu}e_{\mu}-k_{\mu}e_{\nu}\right)G_{n}^{\left(i\right)}\left(p_{i}\right).$
@@
$+\sum_{i}\frac{1}{p_{i}.k}\frac{\mathrm{i}}{4}\varepsilon_{i}^{T}q_{i}\frac{\partial K_{i}\left(-p_{i}\right)}{\partial p_{i\nu}}\frac{\partial\Xi_{i}\left(-p_{i}\right)}{\partial p_{i\mu}}\left(k_{\nu}e_{\mu}-k_{\mu}e_{\nu}\right)G_{n}^{\left(i\right)}\left(p_{i}\right).$ (22)
@@
$\Xi_{i}\left(p_{i}\right)=\mathrm{i}\left(\cancel{p_{i}}+m_{i}\right),$
@@
$\frac{\mathrm{i}}{4}\left(e_{\mu}k_{\nu}-e_{\nu}k_{\mu}\right)q_{i}\varepsilon_{i}^{T}\frac{\partial K_{i}\left(-p_{i}\right)}{\partial p_{i\nu}}\frac{\partial\Xi_{i}\left(-p_{i}\right)}{\partial p_{i\mu}}=$ $-\frac{\mathrm{i}}{4}q_{i}\varepsilon_{i}^{T}\left(e_{\mu}k_{\nu}-e_{\nu}k_{\mu}\right)i\gamma^{\nu}\gamma^{\mu}$
@@
$=$ $-\mathrm{i}q_{i}\varepsilon_{i}^{T}\frac{k_{\mu}e_{\nu}}{p_{i}.k}S_{\left(i\right)}^{\mu\nu}.$
@@
$G_{n+1}\left(e,k;\varepsilon_{i},p_{i}\right)=\sum_{i}\varepsilon_{i}^{T}q_{i}\frac{e.p_{i}}{p_{i}.k}G_{n}^{\left(i\right)}\left(p_{i}\right)-\mathrm{i}\sum_{i}\varepsilon_{i}^{T}q_{i}\frac{e_{\mu}k_{\nu}}{p_{i}.k}\left(\hat{J}_{\left(i\right)}^{\mu\nu}G_{n}^{\left(i\right)}\right)\left(p_{i}\right),$ (23)
@@
$\left(\hat{J}_{\left(i\right)}^{\mu\nu}G_{n}^{\left(i\right)}\right)\left(p_{i}\right)=\mathrm{i}\left\{p_{i}^{\mu}\frac{\partial}{\partial p_{i\nu}}-p_{i}^{\nu}\frac{\partial}{\partial p_{i\mu}}\right\}G_{n}^{\left(i\right)}\left(p_{i}\right)+\frac{\mathrm{i}}{4}\left[\gamma^{\mu},\gamma^{\nu}\right]G_{n}^{\left(i\right)}\left(p_{i}\right).$ (24)
@@
For gluons, the gauge group is $\text{SU}\left(3\right)$ and the Lie algebra has 8 generators. If the external states are quarks, they will be in the fundamental representation of $\text{SU}\left(3\right)$ in which case the generators are given in terms of the 8 Gell-Mann matrices $\lambda_{a}$,
@@
where $\lambda_{i;a}$ are such that the $i$th block is the corresponding Gell-Mann matrix while the other blocks are identity. If the external states are gluons then the generators are in the adjoint representation of $\text{SU}\left(3\right)$,
@@
$G_{n+1}\left(e,k;\varepsilon_{i},p_{i}\right)=\sum_{i}\varepsilon_{i}^{T}g_{i}\frac{e_{a}.p_{i}}{p_{i}.k}T_{i;a}G_{n}^{\left(i\right)}\left(p_{i}\right)-\mathrm{i}\sum_{i}\varepsilon_{i}^{T}g_{i}\frac{e_{a\mu}k_{\nu}}{p_{i}.k}T_{i;a}\left(\hat{J}_{\left(i\right)}^{\mu\nu}G_{n}^{\left(i\right)}\right)\left(p_{i}\right).$ (27)